# Improvise, Adapt, Overcome — Telescopic Adapters for Efficient Fine-tuning of Vision Language Models in Medical Imaging

### v1

@@ 1 Introduction
Medical image segmentation is a dense prediction task that involves labeling each pixel in an image to identify anatomical structures or areas affected by disease [defibation]. With the increasing use of computer-aided diagnosis (CAD) in the last decade [CAD], it has become an important tool in diagnosis [diag], treatment planning [treat3], and accurate patient monitoring [moni].
@@
The success of deep learning in CAD [suc5, suc6] has translated into substantial improvements in medical image segmentation models across diverse imaging modalities, anatomies, and diseases [good]. However, these models are typically trained on specific foreground classes and imaging modalities, limiting their adaptability to new images or additional information sources [general]. Scaling up data for specific segmentation tasks remains challenging due to high costs of data collection and annotation [data_prob], posing challenges for real-world clinical deployment where imaging protocols and diagnostic needs vary widely.
Recent efforts have explored incorporating textual information as auxiliary supervision to overcome these limitations [vlsmadap]. Medical images are frequently accompanied by textual reports containing detailed descriptions of anatomical structures, pathological findings, and spatial relations. These text-based cues complement visual information, particularly when annotations are scarce or image quality is suboptimal [vnlm3]. Vision Language Models (VLMs), such as CLIP [clip], have emerged as powerful foundations for multimodal approaches. CLIP employs contrastive loss on large-scale image-text pairs to align semantically related embeddings in a shared representation space. Building on these capabilities, VLSMs [vlsm2] extend VLMs to dense prediction tasks. CLIPSeg [clipseg] combines frozen CLIP encoders with lightweight transformer-based decoders for open-vocabulary segmentation, while Yu et al. [yu] leverage CLIP with self-supervised mask proposal networks for zero-shot referring image segmentation.
@@ 2.1 Vision Language Models
Vision language models (VLMs) have shown notable progress, enabling cross-modal understanding spanning diverse vision tasks [59, 60]. CLIP [clip], trained on large-scale image–text pairs, has demonstrated strong generalization and zero-shot capabilities [16]. Open source model such as the OpenCLIP implementation [openclip] have further supported large-scale pretraining and reproducibility. Beyond classification, CLIP-based models have been applied to tasks that involve region segmentation and classification through text-driven supervision [18]. These developments have motivated extensions into the medical domain. Models such as PubMedCLIP [14] and MedCLIP [medclip] incorporate biomedical text corpora and improve CLIP alignment in clinical contexts. Other variants [48] introduce self-supervised learning, cross-view consistency, and domain-specific enhancements to improve zero-shot performance in low-resource medical environment. Among these, BiomedCLIP [biomed] demonstrates strong performance in multiorgan retrieval and classification tasks.
Building on the success of CLIP, recent VLSMs aim to take advantage of textual prompts for dense prediction tasks. CLIPSeg [clipseg] enhances CLIP’s architecture by introducing a lightweight transformer decoder that fuses vision and text representations to enable prompt-driven segmentation. Unlike retrieval-based methods, CLIPSeg supports both zero-shot and fine-tuned segmentation scenarios and accommodates ViT backbones. In contrast, Wang et al. [cris] formulates segmentation as a region retrieval problem, employing a CNN-based CLIP vision encoder to align visual features with textual queries. While CRIS [cris] has shown promise in localizing relevant regions, its reliance on a convolutional backbone limits scalability and integration with modern transformer-based pipelines. BiomedCLIPS [biomed] further extends this direction by adapting its encoders for segmentation, though it lacks an E2E pretrained decoder and hence is typically used with task-specific fine-tuning.
@@ 2.2 Parameter Efficient Fine Tuning
Given the prohibitive cost of retraining large-scale models in full, PEFT has emerged as an effective alternative for adapting VLMs to downstream tasks with minimal computational burden. PEFT methods reduce training and memory costs by updating only a small subset of parameters, leaving the majority of pretrained model weights frozen [333]. These methods can be broadly classified into four categories: selective tuning, adapters, prompt-based tuning, and low-rank adaptation. Selective tuning methods restrict optimization to specific components within the architecture. Ben et al. [bitfit] fine-tunes only the bias terms of the attention and feedforward layers, while other approaches apply structured pruning to derive task-specific subnetworks [prune2].
Adapter-based tuning introduces small trainable modules into the frozen backbone to modulate intermediate representations. These modules are typically lightweight bottleneck architectures inserted between transformer sublayers [adaptoer]. Adapter tuning has shown effectiveness in vision language tasks [ex_adap2], enabling modular transfer and multitask learning. Variants such as CLIP-Adapter [ca] adapt pretrained features through learned visual pathways that integrate both adapted and original representations. However, insertion of an adapter increases the number of model parameter inferences, which may lead to higher latency. Recent efforts aim to address this drawback through compression and adapter fusion techniques [fusion].
Prompt tuning replaces manually designed prompts with learnable embeddings. Initially applied to language models, this approach has been extended to vision-language settings by optimizing prompts or visual tokens [prompt1]. These tunable embeddings, when prepended to inputs or injected into intermediate layers, steer the model toward domain-relevant behavior with minimal parameter updates. Prompt tuning is particularly attractive for few-shot and zero-shot settings [fs1].
Low-Rank Adaptation (LoRA) [lora] represents another class of PEFT, where trainable low-rank matrices are injected into linear projections of the frozen model to capture task-specific weight updates. By design, LoRA allows for merging the adapted parameters with the original weights before inference, resulting in no additional latency or memory usage. LoRA assumes task-specific modifications reside in a low-dimensional subspace, enabling efficient fine-tuning. LoRA extension includes rank-adaptive variants [lora2], improved optimization strategies [perform1, perform2], and quantized versions to reduce memory consumption. Wei et al. [imedsam] introduces LoRA to medical image segmentation paradigm for continual learning. Despite its efficiency, LoRA’s performance can be sensitive to the choice of rank and may underperform on tasks requiring more expressive adaptation, especially in high-variance domains like medical imaging.
@@ 3.1 Adapter Placement
Let $\mathcal{A}_{v}$, $\mathcal{A}_{t}$, and $\mathcal{A}_{c}$ denote the adapters in the vision, text, and conditional branches, respectively. Each adapter $\mathcal{A}_{n}$ is inserted in the $n^{th}$ transformer encoder layer $\mathcal{L}_{n}$ of its corresponding modality. Adapters are applied before each residual addition point in $\mathcal{L}_{n}$, resulting in two adapter applications per encoder layer: once after the multi-head self-attention sublayer and once after the feed-forward MLP sublayer as shown in Fig.2 (b). This placement ensures that task-specific modulation occurs before residual addition [adaptoer].
Vision Adapters ($\mathcal{A}_{v}$) are inserted into the first nine transformer layers $\{\mathcal{L}_{1},\mathcal{L}_{2},\ldots,\mathcal{L}_{9}\}$. CLIPSeg utilizes only the intermediate representations of these layers within its decoder, and as such, adapting these specific layers allows the model to progressively adjust the visual features passed to the decoder. Each layer is augmented with two adapter modules placed before the two residual connections, resulting in 18 adapter instances on the vision encoder. This configuration ensures that each encoder block refines its activation before passing it both to the decoder and to the next layer in the stack [clipseg].
@@ 3.2 Adapter Formulation
Building on the formulation introduced in [adaptoer], we design an adapter module that introduces non-linearity, dimensionality compression, and task-specific flexibility into the network without modifying the pretrained backbone as presented Fig.2 (c). The adapted representation is computed as:
@@
We compute $d_{\text{adapter}}$ on a per-layer basis to maintain efficiency and expressiveness. A detailed discussion on this adaptive choice is presented in the next section. The compressed features are then normalized along the feature dimension to ensure stable gradient propagation, followed by a SiLU activation function, which introduces smooth, input-aware non-linearity particularly suited to compact transformation layers [SILU2]. Then a fixed rate dropout (p = 0.1) is applied to regularize the transformation and reduce overfitting.
@@ 3.3 Telescopic Adaptation
Transformer encoders follow a coarse-to-fine representational hierarchy. The early layers predominantly capture low-level or syntactic features, while the deeper layers encode more abstract task-specific semantics [dim]. Imposing large adapters on all layers uniformly not only adds redundancy but may also destabilize training. We address this by dynamically computing the per-layer adapter dimension $d_{\text{adapter}}$ from a base value $d_{\text{base}}=64$, which we empirically select to balance expressiveness and stability. Larger values were found to cause gradient explosion in the early stages of training.
@@ 4.1 Datasets and Evaluation Metrics
To evaluate the effectiveness of our proposed approach, we utilize a diverse collection of medical imaging datasets spanning multiple domains, including both radiology and non-radiology modalities. Following Poudel et al. [vlsm3], we adopt their predefined dataset splits and corresponding text prompts to ensure consistency and comparability in performance benchmarking. Their method generates multiple text prompts per image-mask pair; in our implementation, we randomly select one prompt per sample to construct image-mask-text triplets for evaluation across datasets.
For non-radiology tasks, we utilized three polyp segmentation datasets from endoscopic images: KvasirSEGKvasir-SEG [kvasir], $ClinicDB$ [clinicbd], and $BKAI$ [bakai]. We also included the skin lesion segmentation (ISIC16ISIC-16 [isic16]) dataset for skin cancer detection. For radiology, the breast ultrasound image segmentation dataset ($BUSI$ [busi]) was used.
@@ 4.2 Implementation Details
The training and inference processes for both the baseline and proposed methods were carried out on an NVIDIA RTX A4000 GPU. Mixed-precision training with floating-point 16 was used, employing a batch size of 32. The initial learning rate for the adapter network is set to $1\times 10^{-3}$. A scheduler reduced the learning rate by a factor of 0.3 if no decrease in validation loss was observed over 5 consecutive epochs. Training was stopped early if the validation DSC (%) did not improve in 20 consecutive epochs. All models were optimized using AdamW [adam] with a weight decay of $1\times 10^{-3}$. Each experiment was repeated with three different random seed values to assess consistency and account for variability in prompt sampling. The loss function combined Dice and binary cross-entropy losses as shown in Eq. (6) where scalar values $\lambda_{d}$ and $\lambda_{BCE}$ are set to be 1.5 and 1.
@@ 4.4.1 Ablation Study Over Adapter Placement
Datasets Metric Adapter Configurations • [498k] •• [593k] ••• [613k] Kvasir-SEG [kvasir] DSC (%) 87.35 89.67 89.79 IoU (%) 80.32 83.62 83.50 BKAI [bakai] DSC (%) 85.53 87.09 88.38 IoU (%) 77.77 80.00 81.63 ClinicDB [clinicbd] DSC (%) 85.39 91.28 91.67 IoU (%) 78.45 84.85 85.19 ISIC-16 [isic16] DSC (%) 91.61 92.24 92.18 IoU (%) 85.30 86.16 86.12 BUSI [busi] DSC (%) 70.35 64.33 65.90 IoU (%) 62.45 57.26 59.10
@@
Figure 3: Segmentation results on samples from ClinicDB [clinicbd], ISIC-16 [isic16], BUSI [busi], BKAI [bakai], and Kvasir-SEG [kvasir] datasets (rows 1-5 respectively). Column (a) shows the original medical images, followed by segmentation masks from: (b) ZSS (zero-shot segmentation using CLIPSeg [clipseg]), (c) CLIPSeg [clipseg] (E2E fine-tuned), (d) SAN [san], (e) CRIS [cris] (E2E fine-tuned), (f) CLIPSeg SA VLC [vlsmadap], (g) CLIPSeg DA VLC [vlsmadap], (h) I-MedSAM [imedsam], (i) Telescopic Adapters, and (j) GT (ground truth).
@@
The addition of text adapters (••) to the vision branch baseline (•) provides a substantial performance uplift across most datasets, underscoring the importance of textual context. For example, on the ClinicDB [clinicbd] dataset, this addition improves the DSC from 85.39% to 91.28% and the IoU from 78.45% to 84.85%.
Our proposed model (•••), which includes the conditional branch, consistently delivers the best or most competitive performance. This final component adds further refinement for a minimal parameter overhead of only 20k, achieving the highest scores for both metrics on the BKAI [bakai] and ClinicDB datasets. Which confirms that the conditional adapters effectively harmonize the features from the vision and text branches.
Notably, for the BUSI [busi] dataset, the vision branch baseline (•) yields the best results. This suggests that for certain datasets with less complex inter-modal dependencies, the additional branches may not provide a benefit. However, the superior performance of the three (i.e. Vision, Text, and Conditional) branches model on approximately four out of five diverse datasets validates it as the most robust and effective configuration, offering an optimal balance between parameter efficiency and high generalization capability.
@@
CLIPSeg [clipseg]
150M
CLIPSeg [clipseg]
150M
CRIS [cris]
147M
I-MedSAM [imedsam]
1.6M
SAN [san]
8.4M
CLIPSeg DA VLC [vlsmadap]
3.2M
CLIPSeg SA VLC [vlsmadap]
4.2M
Telescopic Adapters
[OURS] 613k
Kvasir-SEG [kvasir] DSC (%) 53.42 87.69 89.43 86.36 69.58 89.06 86.80 89.79
@@
Table 4: Quantitative benchmark of segmentation performance across five medical datasets: Kvasir-SEG [kvasir], BKAI [bakai], ClinicDB [clinicbd], ISIC-16 [isic16], and BUSI [busi]. We compare our Telescopic Adapters (OURS) against: (i) the Zero-Shot (ZS) baseline (CLIPSeg [clipseg]), (ii) End-to-End (E2E) fine-tuning (CLIPSeg [clipseg], CRIS [cris]), (iii) LoRA-based tuning (I-MedSAM [imedsam]), and (iv) other adapter-based methods (SAN [san], CLIPSeg DA/SA VLC [vlsmadap]). Performance is reported in DSC (%) and IoU (%). Parameter counts shown for PEFT methods and End-to-End fine-tuning are trainable; for Zero-Shot segmentation, they are the total model parameters.

### v2

@@ 1 Introduction
Medical image segmentation is a dense prediction task that involves labeling each pixel in an image to identify anatomical structures or areas affected by disease [41]. With the increasing use of computer-aided diagnosis (CAD) in the last decade [9], it has become an important tool in diagnosis [11], treatment planning [15], and accurate patient monitoring [12].
@@
The success of deep learning in CAD [50, 56] has translated into substantial improvements in medical image segmentation models across diverse imaging modalities, anatomies, and diseases [31]. However, these models are typically trained on specific foreground classes and imaging modalities, limiting their adaptability to new images or additional information sources [5]. Scaling up data for specific segmentation tasks remains challenging due to high costs of data collection and annotation [16], posing challenges for real-world clinical deployment where imaging protocols and diagnostic needs vary widely.
Recent efforts have explored incorporating textual information as auxiliary supervision to overcome these limitations [6]. Medical images are frequently accompanied by textual reports containing detailed descriptions of anatomical structures, pathological findings, and spatial relations. These text-based cues complement visual information, particularly when annotations are scarce or image quality is suboptimal [28]. Vision Language Models (VLMs), such as CLIP [39], have emerged as powerful foundations for multimodal approaches. CLIP employs contrastive loss on large-scale image-text pairs to align semantically related embeddings in a shared representation space. Building on these capabilities, VLSMs [14] extend VLMs to dense prediction tasks. CLIPSeg [32] combines frozen CLIP encoders with lightweight transformer-based decoders for open-vocabulary segmentation, while Yu et al. [48] leverage CLIP with self-supervised mask proposal networks for zero-shot referring image segmentation.
@@ 2.1 Vision Language Models
Vision language models (VLMs) have shown notable progress, enabling cross-modal understanding spanning diverse vision tasks [47, 49]. CLIP [39], trained on large-scale image–text pairs, has demonstrated strong generalization and zero-shot capabilities [17]. Open source model such as the OpenCLIP implementation [24] have further supported large-scale pretraining and reproducibility. Beyond classification, CLIP-based models have been applied to tasks that involve region segmentation and classification through text-driven supervision [18]. These developments have motivated extensions into the medical domain. Models such as PubMedCLIP [10] and MedCLIP [44] incorporate biomedical text corpora and improve CLIP alignment in clinical contexts. Other variants [43] introduce self-supervised learning, cross-view consistency, and domain-specific enhancements to improve zero-shot performance in low-resource medical environment. Among these, BiomedCLIP [52] demonstrates strong performance in multiorgan retrieval and classification tasks.
Building on the success of CLIP, recent VLSMs aim to take advantage of textual prompts for dense prediction tasks. CLIPSeg [32] enhances CLIP’s architecture by introducing a lightweight transformer decoder that fuses vision and text representations to enable prompt-driven segmentation. Unlike retrieval-based methods, CLIPSeg supports both zero-shot and fine-tuned segmentation scenarios and accommodates ViT backbones. In contrast, Wang et al. [42] formulates segmentation as a region retrieval problem, employing a CNN-based CLIP vision encoder to align visual features with textual queries. While CRIS [42] has shown promise in localizing relevant regions, its reliance on a convolutional backbone limits scalability and integration with modern transformer-based pipelines. BiomedCLIPS [52] further extends this direction by adapting its encoders for segmentation, though it lacks an E2E pretrained decoder and hence is typically used with task-specific fine-tuning.
@@ 2.2 Parameter Efficient Fine Tuning
Given the prohibitive cost of retraining large-scale models in full, PEFT has emerged as an effective alternative for adapting VLMs to downstream tasks with minimal computational burden. PEFT methods reduce training and memory costs by updating only a small subset of parameters, leaving the majority of pretrained model weights frozen [29]. These methods can be broadly classified into four categories: selective tuning, adapters, prompt-based tuning, and low-rank adaptation. Selective tuning methods restrict optimization to specific components within the architecture. Ben et al. [2] fine-tunes only the bias terms of the attention and feedforward layers, while other approaches apply structured pruning to derive task-specific subnetworks [21].
Adapter-based tuning introduces small trainable modules into the frozen backbone to modulate intermediate representations. These modules are typically lightweight bottleneck architectures inserted between transformer sublayers [22]. Adapter tuning has shown effectiveness in vision language tasks [7], enabling modular transfer and multitask learning. Variants such as CLIP-Adapter [13] adapt pretrained features through learned visual pathways that integrate both adapted and original representations. However, insertion of an adapter increases the number of model parameter inferences, which may lead to higher latency. Recent efforts aim to address this drawback through compression and adapter fusion techniques [37].
Prompt tuning replaces manually designed prompts with learnable embeddings. Initially applied to language models, this approach has been extended to vision-language settings by optimizing prompts or visual tokens [26]. These tunable embeddings, when prepended to inputs or injected into intermediate layers, steer the model toward domain-relevant behavior with minimal parameter updates. Prompt tuning is particularly attractive for few-shot and zero-shot settings [33].
Low-Rank Adaptation (LoRA) [23] represents another class of PEFT, where trainable low-rank matrices are injected into linear projections of the frozen model to capture task-specific weight updates. By design, LoRA allows for merging the adapted parameters with the original weights before inference, resulting in no additional latency or memory usage. LoRA assumes task-specific modifications reside in a low-dimensional subspace, enabling efficient fine-tuning. LoRA extension includes rank-adaptive variants [51], improved optimization strategies [27, 4], and quantized versions to reduce memory consumption. Wei et al. [45] introduces LoRA to medical image segmentation paradigm for continual learning. Despite its efficiency, LoRA’s performance can be sensitive to the choice of rank and may underperform on tasks requiring more expressive adaptation, especially in high-variance domains like medical imaging.
@@ 3.1 Adapter Placement
Let $\mathcal{A}_{v}$, $\mathcal{A}_{t}$, and $\mathcal{A}_{c}$ denote the adapters in the vision, text, and conditional branches, respectively. Each adapter $\mathcal{A}_{n}$ is inserted in the $n^{th}$ transformer encoder layer $\mathcal{L}_{n}$ of its corresponding modality. Adapters are applied before each residual addition point in $\mathcal{L}_{n}$, resulting in two adapter applications per encoder layer: once after the multi-head self-attention sublayer and once after the feed-forward MLP sublayer as shown in Fig.2 (b). This placement ensures that task-specific modulation occurs before residual addition [22].
Vision Adapters ($\mathcal{A}_{v}$) are inserted into the first nine transformer layers $\{\mathcal{L}_{1},\mathcal{L}_{2},\ldots,\mathcal{L}_{9}\}$. CLIPSeg utilizes only the intermediate representations of these layers within its decoder, and as such, adapting these specific layers allows the model to progressively adjust the visual features passed to the decoder. Each layer is augmented with two adapter modules placed before the two residual connections, resulting in 18 adapter instances on the vision encoder. This configuration ensures that each encoder block refines its activation before passing it both to the decoder and to the next layer in the stack [32].
@@ 3.2 Adapter Formulation
Building on the formulation introduced in [22], we design an adapter module that introduces non-linearity, dimensionality compression, and task-specific flexibility into the network without modifying the pretrained backbone as presented Fig.2 (c). The adapted representation is computed as:
@@
We compute $d_{\text{adapter}}$ on a per-layer basis to maintain efficiency and expressiveness. A detailed discussion on this adaptive choice is presented in the next section. The compressed features are then normalized along the feature dimension to ensure stable gradient propagation, followed by a SiLU activation function, which introduces smooth, input-aware non-linearity particularly suited to compact transformation layers [53]. Then a fixed rate dropout (p = 0.1) is applied to regularize the transformation and reduce overfitting.
@@ 3.3 Telescopic Adaptation
Transformer encoders follow a coarse-to-fine representational hierarchy. The early layers predominantly capture low-level or syntactic features, while the deeper layers encode more abstract task-specific semantics [35]. Imposing large adapters on all layers uniformly not only adds redundancy but may also destabilize training. We address this by dynamically computing the per-layer adapter dimension $d_{\text{adapter}}$ from a base value $d_{\text{base}}=64$, which we empirically select to balance expressiveness and stability. Larger values were found to cause gradient explosion in the early stages of training.
@@ 4.1 Datasets and Evaluation Metrics
To evaluate the effectiveness of our proposed approach, we utilize a diverse collection of medical imaging datasets spanning multiple domains, including both radiology and non-radiology modalities. Following Poudel et al. [38], we adopt their predefined dataset splits and corresponding text prompts to ensure consistency and comparability in performance benchmarking. Their method generates multiple text prompts per image-mask pair; in our implementation, we randomly select one prompt per sample to construct image-mask-text triplets for evaluation across datasets.
For non-radiology tasks, we utilized three polyp segmentation datasets from endoscopic images: KvasirSEGKvasir-SEG [25], $ClinicDB$ [3], and $BKAI$ [34]. We also included the skin lesion segmentation (ISIC16ISIC-16 [19]) dataset for skin cancer detection. For radiology, the breast ultrasound image segmentation dataset ($BUSI$ [1]) was used.
@@ 4.2 Implementation Details
The training and inference processes for both the baseline and proposed methods were carried out on an NVIDIA RTX A4000 GPU. Mixed-precision training with floating-point 16 was used, employing a batch size of 32. The initial learning rate for the adapter network is set to $1\times 10^{-3}$. A scheduler reduced the learning rate by a factor of 0.3 if no decrease in validation loss was observed over 5 consecutive epochs. Training was stopped early if the validation DSC (%) did not improve in 20 consecutive epochs. All models were optimized using AdamW [30] with a weight decay of $1\times 10^{-3}$. Each experiment was repeated with three different random seed values to assess consistency and account for variability in prompt sampling. The loss function combined Dice and binary cross-entropy losses as shown in Eq. (6) where scalar values $\lambda_{d}$ and $\lambda_{BCE}$ are set to be 1.5 and 1.
@@ 4.4.1 Ablation Study Over Adapter Placement
Datasets Metric Adapter Configurations • [498k] •• [593k] ••• [613k] Kvasir-SEG [25] DSC (%) 87.35 89.67 89.79 IoU (%) 80.32 83.62 83.50 BKAI [34] DSC (%) 85.53 87.09 88.38 IoU (%) 77.77 80.00 81.63 ClinicDB [3] DSC (%) 85.39 91.28 91.67 IoU (%) 78.45 84.85 85.19 ISIC-16 [19] DSC (%) 91.61 92.24 92.18 IoU (%) 85.30 86.16 86.12 BUSI [1] DSC (%) 70.35 64.33 65.90 IoU (%) 62.45 57.26 59.10
@@
Figure 3: Segmentation results on samples from ClinicDB [3], ISIC-16 [19], BUSI [1], BKAI [34], and Kvasir-SEG [25] datasets (rows 1-5 respectively). Column (a) shows the original medical images, followed by segmentation masks from: (b) ZSS (zero-shot segmentation using CLIPSeg [32]), (c) CLIPSeg [32] (E2E fine-tuned), (d) SAN [46], (e) CRIS [42] (E2E fine-tuned), (f) CLIPSeg SA VLC [6], (g) CLIPSeg DA VLC [6], (h) I-MedSAM [45], (i) Telescopic Adapters, and (j) GT (ground truth).
@@
The addition of text adapters (••) to the vision branch baseline (•) provides a substantial performance uplift across most datasets, underscoring the importance of textual context. For example, on the ClinicDB [3] dataset, this addition improves the DSC from 85.39% to 91.28% and the IoU from 78.45% to 84.85%.
Our proposed model (•••), which includes the conditional branch, consistently delivers the best or most competitive performance. This final component adds further refinement for a minimal parameter overhead of only 20k, achieving the highest scores for both metrics on the BKAI [34] and ClinicDB datasets. Which confirms that the conditional adapters effectively harmonize the features from the vision and text branches.
Notably, for the BUSI [1] dataset, the vision branch baseline (•) yields the best results. This suggests that for certain datasets with less complex inter-modal dependencies, the additional branches may not provide a benefit. However, the superior performance of the three (i.e. Vision, Text, and Conditional) branches model on approximately four out of five diverse datasets validates it as the most robust and effective configuration, offering an optimal balance between parameter efficiency and high generalization capability.
@@
CLIPSeg [32] 150M CLIPSeg [32] 150M CRIS [42] 147M I-MedSAM [45] 1.6M SAN [46] 8.4M CLIPSeg DA VLC [6] 3.2M CLIPSeg SA VLC [6] 4.2M Telescopic Adapters [OURS] 613k
Kvasir-SEG [25] DSC (%) 53.42 87.69 89.43 86.36 69.58 89.06 86.80 89.79
@@
Table 4: Quantitative benchmark of segmentation performance across five medical datasets: Kvasir-SEG [25], BKAI [34], ClinicDB [3], ISIC-16 [19], and BUSI [1]. We compare our Telescopic Adapters (OURS) against: (i) the Zero-Shot (ZS) baseline (CLIPSeg [32]), (ii) End-to-End (E2E) fine-tuning (CLIPSeg [32], CRIS [42]), (iii) LoRA-based tuning (I-MedSAM [45]), and (iv) other adapter-based methods (SAN [46], CLIPSeg DA/SA VLC [6]). Performance is reported in DSC (%) and IoU (%). Parameter counts shown for PEFT methods and End-to-End fine-tuning are trainable; for Zero-Shot segmentation, they are the total model parameters.